OpenStudy (anonymous):

What are the solutions of 4x2 + x = -3? My answer: -1 + 7i over 8

hartnn (hartnn):

b^2-4ac = 1-4*4*3 = 1-48 = -47

hartnn (hartnn):

i think u got -49

OpenStudy (anonymous):